OpenStudy (anonymous):

If a ≠ b and 1/x + 1/a = 1/b , then x =

OpenStudy (dumbcow):

combine LHS \[\frac{a +x}{ax} = \frac{1}{b}\] cross-multiply \[ab + bx = ax\] get "x" terms on same side, factor out the "x" \[x(a-b) = ab\] divide \[x = \frac{ab}{a-b}\]